OpenStudy (anonymous):

find the horizontal asymtote of the equation y=x-2/16-x^2

OpenStudy (anonymous):

Do you mean \[\frac{x-2}{16}\] or \[x-\frac{2}{16}\]

OpenStudy (anonymous):

Is it \[y=\frac{x-2}{16-x^2}\]

OpenStudy (anonymous):

\[y=\frac{ x-2 }{ 16-x^2 }\]

OpenStudy (anonymous):

i think you gave the intercept of the the equation

OpenStudy (anonymous):

Divide numerator and denominator by \[x^2\] Then, you apply limit for \[\lim_{x\rightarrow \pm \infty}\]

OpenStudy (anonymous):

horizontal asymptote means, the value that y-approaches to as x becomes +ve or -ve infinite. steps: 1. make the substitution x=1/u 2. simplify 3. put u=0 (as x->±inf, u->0) 4. you will get a value of y

OpenStudy (anonymous):

i think the u in there should be the lowest degree of x isn't it?

OpenStudy (anonymous):

You end up getting same answer, but yeah, in general, you use lower degree one.

OpenStudy (anonymous):

ok i'll try it thanks

OpenStudy (anonymous):

no need. just make the substitution and it will always work.

OpenStudy (anonymous):

there is a shortcut too.... (dont tell your prof ;)) 1. find highest power of x = n (in this case, 2) 2. y = ratio of the coefficients of the x^x terms on top and bottom in this case. y = 0/(-1) = 0 so, the horiz asymptote is y=0

OpenStudy (anonymous):

ahaha thanks i'll keep it for myself^_^